OpenStudy (anonymous):

What is the volume of a cube with a side length of 1/11 of a unit?

OpenStudy (anonymous):

length x width x high

OpenStudy (anonymous):

\[\left(\frac{1}{11}\right)^3=\frac{1}{1331} \]